# Entanglement measures of Majorana bound states

### v1

@@ Abstract
Majorana bound states emerge in topological superconductors as zero-energy edge states exhibiting spatial nonlocality. Despite the enormous advances, the detection of Majorana bound states is still challenging mainly because topologically trivial Andreev bound states produce similar signatures. In this work we consider a topological superconductor with Majorana bound states coupled to quantum dots and investigate the dynamics of their quantum correlations with the aim to explore their entanglement properties. In particular, we characterize entanglement by using concurrence and discord, which are also complemented by the entanglement dynamics and return probability. We find that Majorana bound states at truly zero energy can transform an initially entangled system into its classical state, while they can create maximally entangled states at a finite energy overlap. Interestingly, we show that the system can generate a maximally entangled state between MBSs and a quantum dot by simply controlling the Majorana nonlocality. We demonstrate that these results hold in the scenarios when the initial state is either maximally entangled or separable, albeit in the latter maximally entangled states are achieved in the long time dynamics. Furthermore, we contrast our findings with those produced by a regular fermion and obtain very distinct entanglement signatures. Our work offers an alternative approach to characterize Majorana bound states, which can be also useful towards their utilization for quantum information tasks.
@@ I Introduction
Majorana bound states (MBSs) have become one of the central topics in condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8, 9] largely due to their promising properties for fault tolerant quantum computation [10, 11, 12, 13, 14]. MBSs were initially predicted to appear in the topological phase of spinless $p$-wave superconductors, which later were shown to be realized by combining spin-orbit coupling, an external magnetic field, and conventional spin-singlet $s$-wave superconductivity, see e.g., Ref. [9]. MBSs emerge at zero-energy with their wavefunctions located at the system edges, thus exhibiting an inherent spatial nonlocality. In this regard, truly zero-energy MBSs are needed for realizing qubits that are robust against local perturbations [10]. However, zero-energy signatures can be also produced by trivial Andreev bound states [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], which proliferate in real devices and have challenged the detection of MBSs [7].
@@
A less explored property of MBSs is their spatial nonlocality, which can be revealed by inducing a finite spatial overlap between Majorana wavefunctions [18, 29, 30, 31]. In this case, MBSs acquire a finite energy splitting, which, although not beneficial for realizing topological qubits, is useful for distinguishing the inherent Majorana nature. Following this idea, it has been recently shown that testing the Majorana nonlocality allows to distinguish between MBSs and trivial Andreev bound states [32, 33, 34, 35, 36, 37, 38, 39, 40]. Even though accessing the Majorana nonlocality might be difficult, it offers a solid way for the unambiguous detection of MBSs [34]. Moreover, since the spatial nonlocality reflects the MBSs at spatially separated regions, it is natural to wonder if such a spatial nonlocality carries quantum correlations and influences entanglement properties of the system.
@@
In this work we consider a topological superconductor with two MBSs coupled to quantum dots (QDs) as in Fig. 1 and investigate their entanglement properties. In particular, we characterize entanglement by exploring the dynamics of concurrence and discord, quantum correlations that we also complement by obtaining the entanglement dynamics and return probability. We find that zero-energy MBSs, which are completely localized at the edges, can evolve an initial state of maximally entangled QDs to a classical state of the system. Interestingly, we discover that a finite Majorana overlap enables maximally entangled states between MBSs and a QD, an effect that is fully controllable by the couplings to the QDs. We show that the entanglement generation occurs when the initial state is maximally entangled and, surprisingly, also for an initial separable state, where maximal entanglement in the latter is achieved in the long time dynamics. In both cases, the entanglement signatures are intimmely tied to energy overlap between MBSs and, therefore, associated to their spatial nonlocality. We also show that QDs mediated by a regular fermion does not produce the entanglement signatures found for MBSs.
@@
Figure 1: Sketch of the studied system: a topological superconductor (red) hosting MBSs $\gamma_{1,2}$ is coupled by $\lambda_{1,2}$ to two QDs (blue) with onsite energies $\varepsilon_{1,2}$. The spatial overlap between $\gamma_{1}$ and $\gamma_{2}$ originates and energy splitting represented by $\varepsilon_{\rm M}$.
@@
The paper is organised as follows: In Section II, we describe the model Hamiltonians of the systems and obtain their eigenspectrums. In Sec. III, we briefly describe the different measures of quantum correlations and state dynamical quantities. Following that, in Sec. IV, we compute the state dynamical functions and quantum correlation measures and give analytical expressions for all quantities for different initial states. Subsequently, we discuss the results in the later part of the section. Finally, in Section V, we summarise and conclude the results.
@@ II The Model Hamiltonians and methods
We are interested in exploring quantum correlations and entanglement signatures of MBSs. For this purpose, we consider a topological superconductor hosting MBSs and couple them to two QDs, as shown in Fig. 1. Moreover, we contrast this topological model by replacing the nonlocal fermion, arising due to both MBSs, with a regular fermion. In what follows we present these two models, discuss their properties, and introduce the methods used in this work to quantify entanglement measures.
@@ II.1 QDs coupled through MBSs
where $\gamma_{i}$ represents a Majorana operator, $\epsilon_{M}$ is the energy of their splitting between MBSs, $d_{j}^{\dagger}(d_{j})$ creates (annihilates) a fermionic state in the QDs, and $\lambda_{j}$ characterizes the coupling between QDs MBSs. We note that the energy splitting $\epsilon_{M}$ reflects the Majorana nonlocality and can be approximated as $\epsilon_{M}\sim e^{-L/l}$, where L is the length of the topological superconductor and $l$ is the Majorana localisation length.
To access quantum correlations and entanglement, it is convenient to write Eq. (1) in terms of fermionic operators and then handle it in a number represent basis. Thus, the two Majorana operators are transformed into a nonlocal fermion by using $\gamma_{1}=i(f^{\dagger}-f)$, $\gamma_{2}=f^{\dagger}+f$, where $f$ represents the nonlocal fermion made of Majorana operators. Now, the two QDs plus the nonlocal fermion represent a three fermion system, with number operators $n_{d_{i}}=d^{\dagger}_{i}d_{i}$ for the QDs while $n_{f}=f^{\dagger}f$ for the nonlocal fermion operator. This allows us to define a three qubit system which will be analyzed next.
Before going further, we adopt some simplifications that will allow us to derive analytical expressions. The on-site energies of the QDs are set to zero $\epsilon_{d_{1}}=\epsilon_{d_{2}}=0$, the coupling to the left QD is set to unity $\lambda_{1}=1$, $\omega=\epsilon_{M}/2$, $\lambda_{2}=\lambda$, and $\hbar=1$. Hence, tuning $\lambda$ and $\omega$ allow us to control the properties of the entire system. We now proceed to write Eq. (1) in the number representation basis defined by |$\ket{n_{d_{1}},n_{d_{2}},n_{f}}$. In this regard, we note that, since the Hamiltonian commutes with the fermion parity operator $(-1)^{N}$, with $N$ being the total number of fermions, the odd and the even sectors do not mix and can be solved separately. We focus on the even sector with basis states |$\ket{000}$, |$\ket{011}$, |$\ket{101}$, and |$\ket{110}$. The even Hamiltonian can be then written as,
@@
where |$\ket{E_{1}}=(1,(\omega-\Delta)/(\lambda-1),(\omega-\Delta)/(\lambda-1),1)^{T}$ and T is the transpose operation. Moreover, |$\ket{E_{2}}=\ket{E_{1}(\delta\rightarrow-\delta)}$, |$\ket{E_{3(4)}}=\ket{E_{1(2)}(\Delta\rightarrow\Delta_{1})}$. We note that all the eigenvalues in Eqs. (3) have contributions from both $\lambda$ and $\omega$, while the respective eigenstates in Eq. (4) exhibit contributions from all configurations the even ket states listed above Eq. (2). This distribution of the Hamiltonian parameters in the eigenvalues and eigenvectors arises precisely from the nonlocal fermionic structure of MBSs.
@@ II.2 QDs coupled through a normal fermion
where $\Delta_{\pm}=\omega\pm\sqrt{1+\lambda^{2}+\omega^{2}}$. Moreover, the associated unnormalized eigenstates for increasing values of $i$ in $E^{\prime}_{i}$ are given by
@@
II.3 Methods to quantify entanglement and the state dynamics
@@
where $\rho$ is the density matrix of the composite three qubit system, ${\rm Tr}_{d^{\prime}}$ represents the partial trace operation over a subsystem such that $\rho_{d}$ describes a two qubit system. We remind that the three qubit system is formed by the two QDs and either the nonlocal fermion due to MBSs or a normal fermion, see Section II for more details on the models. The density matrix in Eq. (8) is obtained by using a standard approach $\rho=\ket{\psi}\bra{\psi}$ |, where |$\ket{\psi}$ corresponds to the state of the system [41]. We anticipate that, because we are interested in the dynamical properties, we use the time-evolved states of the Majorana and normal fermion systems discussed in previous section, taking into account different initial states in both systems. Then, all the bipartite quantum correlations can be studied by calculating the reduced density matrix $\rho_{d}$ given by Eq. (8), thus providing a starting point for exploring various entanglement measures. In particular, we will address the concurrence and discord because they provide a unify way to quantify entanglement and quantum correlations [41]. To complement these quantities, we will also address the state dynamics by employing the return probability and the entanglement dynamics. To make this work self contained, in what follows we briefly describe the fundamental aspects of these quantities.
@@ II.3.1 Concurrence
$C=\max(\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{1},0)\,,$ (9)
which correspond to the maximum of the eigenvalues $\lambda_{i}$, with them being the eigenvalues in decreasing order of the matrix $R=\sqrt{\sqrt{\rho_{d}}\tilde{\rho_{d}}\sqrt{\rho_{d}}}$. The matrix $\tilde{\rho_{d}}$ is defined as $\tilde{\rho_{d}}=\sigma_{y}\otimes\sigma_{y}\rho_{d}^{*}\sigma_{y}\otimes\sigma_{y}$, where $\rho_{d}^{*}$ is the complex conjugate reduced density matrix $\rho_{d}$ given by Eq. (8). In our systems, the reduced density matrices in the basis states of |$\ket{00}$, |$\ket{01}$, |$\ket{10}$, and |$\ket{11}$ acquire the following form
@@
$C=2\max(|x|-\sqrt{uv},|y|-\sqrt{w_{1}w_{2}},0)\,.$ (11)
Before going further, a few comments are in order at this stage. The concurrence given by Eq. (11) remains finite if the off-diagonal correlation functions ($x,y$) dominate over their diagonal counterparts ($\sqrt{uv},\sqrt{w_{1}w_{2}}$); otherwise, $C$ goes to zero. A zero concurrence $C=0$ signals that appearance of an unentangled state of the system but, importantly, this does not imply that the total quantum correlations in the system are completely lost or absent [44]. The approach discussed here will be used later to obtain the concurrence and investigate entanglement in a bipartite system.
@@ II.3.2 Quantum discord
To quantify quantum correlations beyond entanglement, we focus on the quantum discord because it measures purely quantum correlations in a bipartite system. The quantum discord is of particular relevance when exploring the dynamics of the system, which, although acquiring a non-entangled state at certain times, may still exhibit finite quantum correlations [45, 46, 47, 48, 49, 50, 51, 52]. The discord is a bipartite correlation measure, and, in this sense, it is similar to concurrence. It is defined by the difference between two classically equivalent expressions of the total correlation, characterized by the quantum mutual information [53]. It is thus useful to first introduce the quantum mutual information, which, for subsystems A and B, is characterized by [53]
@@
where $S(\rho)$ is the entropy associated to $\rho$ and calculated as $S(\rho)=-Tr(\rho\log_{2}\rho)$, while $C_{\theta,\phi}(\rho_{A|B})$ is the conditional entropy of A given the state of B. Moreover, $\rho_{A(B)}$ is the reduced density matrix of A(B), $\rho_{AB}$ is the composite reduced density matrix of the subsystem AB, and $\rho_{A|B}$ represents the conditional density matrix. Furthermore, the conditional entropy is obtained as [53]
@@
where $\theta$ and $\phi$ are angles that parametrize the measurement basis of B, $\{B_{\tilde{\kappa}}\}=\{\ket{\tilde{\kappa}}\bra{\tilde{\kappa}}\}$ | } is a complete set of projection operators corresponding to local measurements on B, and min indicates that a minimization operation with respect to $\{B_{\tilde{\kappa}}\}$ is carried out to find $C$. Also, $p_{\tilde{\kappa}}$ on the right hand side of Eq. (13) denotes the probability of measurement outcome $\tilde{\kappa}$ and is defined as $p_{\tilde{\kappa}}=\text{Tr}[(\sigma_{0}\otimes B_{\tilde{\kappa}})\rho_{AB}(\sigma_{0}\otimes B_{\tilde{\kappa}})]$, with $\sigma_{0}$ representing the identity matrix for A. Moreover, $\rho_{\tilde{\kappa}}=(\sigma_{0}\otimes B_{\tilde{\kappa}})\rho_{AB}(\sigma_{0}\otimes B_{\tilde{\kappa}})/p_{\tilde{\kappa}}$, which represents that it is the density matrix $\rho_{AB}$ conditioned to the measurement outcomes $\tilde{\kappa}$ and weighted over different outcomes of marginal conditional entropies $S(\rho_{A|B_{\tilde{\kappa}}})$. The term $\rho_{A|B_{\tilde{\kappa}}}$ represents the conditional density matrices in the $|\tilde{\kappa}\rangle$ basis. In terms of all the above considerations, the quantum discord is defined as [53]
$D_{AB}={\rm min}[I(\rho_{AB})-J(\rho_{AB})]\,.$ (14)
where $I(\rho_{AB})$ and $J(\rho_{AB})$ are given by Eqs. (12). The minimization of the conditional entropy makes the quantum discord a quantifier of truly quantum correlation. In Appendix A, we provide step-by-step calculations of discord in different subsections for both MS and NFS.

### v2

@@ Abstract
Majorana bound states emerge in topological superconductors as zero-energy edge states exhibiting spatial nonlocality. Despite the enormous advances, the detection of Majorana bound states is still challenging, mainly because topologically trivial Andreev bound states produce similar signatures. In this work, we consider a topological superconductor with Majorana bound states coupled to quantum dots and investigate the dynamics of their quantum correlations with the aim to explore their entanglement properties. In particular, we characterize entanglement by using concurrence and discord, which are also complemented by entanglement dynamics and return probability. We find that Majorana bound states at truly zero energy can transform an initially entangled system into its classical state, while they can create maximally entangled states at a finite energy overlap. Interestingly, we show that the system can generate a maximally entangled state between MBSs and a quantum dot by simply controlling the Majorana nonlocality. We demonstrate that these results hold in scenarios when the initial state is either maximally entangled or separable, albeit in the latter, maximally entangled states are achieved in the long time dynamics. Furthermore, we contrast our findings with those produced by a regular fermion and obtain very distinct entanglement signatures. Our work offers an alternative approach to characterizing Majorana bound states, which can also be useful towards their utilization for quantum information tasks.
@@ I Introduction
Majorana bound states (MBSs) have become one of the central topics in condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8, 9] largely due to their promising properties for fault tolerant quantum computation [10, 11, 12, 13, 14]. MBSs were initially predicted to appear in the topological phase of spinless $p$-wave superconductors, which later were shown to be realized by combining spin-orbit coupling, an external magnetic field, and conventional spin-singlet $s$-wave superconductivity; see, e.g., Ref. [9]. MBSs emerge at zero energy with their wavefunctions located at the system edges, thus exhibiting an inherent spatial nonlocality. In this regard, truly zero-energy MBSs are needed for realizing qubits that are robust against local perturbations [10]. Moreover, zero-energy signatures can be also produced by trivial Andreev bound states (ABSs) [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], which proliferate in real devices and have challenged the detection of MBSs [7].
@@
A less explored property of MBSs is their spatial nonlocality, which can be revealed by inducing a finite spatial overlap between Majorana wavefunctions [18, 30, 31, 32]. In this case, MBSs acquire finite energy splitting, which, although not beneficial for realizing topological qubits, is useful for distinguishing the inherent Majorana nature. Following this idea, it has been recently shown that testing the Majorana nonlocality allows to distinguish between MBSs and trivial Andreev bound states [33, 34, 35, 36, 37, 38, 39, 40, 41]. Majorana nonlocality has also been explored by interference effects when coupling topological superconductors and quantum dots (QDs) [42, 43, 35], with shot noise signatures signaling the nonlocal nature of MBSs [44]. This interference effect has also been shown to provide a way to distinguish between MBSs and Andreev bound states by exploiting the Josephson effect [39, 40], nonlocal conductance [39, 45], and also finite frequency noise [46]; see also Refs. [47, 48, 49, 50, 51] for noise signatures of MBSs. Furthermore, the Majorana nonlocality can be detected by the Majorana fractional entropy [52, 53, 54, 55], which can be of particular relevance for experimental measurements of entropy in nanosystems [56, 57, 58, 59, 60, 61]. The Majorana nonlocaly thus offers a solid way for the unambiguous detection of MBSs, and, despite its challenging characterization, there already exist experimental evidence supporting its utility [35]. Moreover, since the spatial nonlocality reflects the MBSs at spatially separated regions, it is natural to wonder if such a spatial nonlocality carries quantum correlations and influences the entanglement properties of the system [62].
@@
In this work, we consider a topological superconductor with two MBSs coupled to QDs, as shown in Fig. 1. We investigate the entanglement properties of MBSs. In particular, we characterize entanglement by exploring the dynamics of concurrence and discord, quantum correlations that we also complement by obtaining the entanglement dynamics and return probability. We find that zero-energy MBSs, which are completely localized at the edges, can evolve an initial state of maximally entangled QDs to a classical state of the system. Interestingly, we discover that a finite Majorana overlap enables maximally entangled states between MBSs and a QD, an effect that is fully controllable by the couplings to the QDs. We show that entanglement generation occurs when the initial state is maximally entangled and, surprisingly, also for an initial separable state, where maximal entanglement in the latter is achieved in the long time dynamics. In both cases, the entanglement signatures are intimately tied to energy overlap between MBSs and, therefore, associated with their spatial nonlocality. We also show that QDs mediated by a regular fermion do not produce the entanglement signatures found for MBSs. While entanglement generation has been explored through other mechanisms—such as due to photons [63, 64, 65] and cavities [66, 67], our work exploits the spatial nonlocality of MBSs, which then provides an alternative way for characterizing MBSs and highlights their potential for quantum information. Given the current experimental advances on quantum correlations, our findings can be measured by performing quantum state tomography [68, 69] which allows to reconstruct the system’s density matrix.
@@
Figure 1: Sketch of the system: a topological superconductor (red) hosting MBSs $\gamma_{1,2}$ is coupled by $\lambda_{1,2}$ to two QDs (blue) with onsite energies $\varepsilon_{1,2}$. The energy associated with the spatial overlap between $\gamma_{1}$ and $\gamma_{2}$ is represented by $\varepsilon_{\rm M}$.
@@
The paper is organised as follows: In Section II, we describe the model Hamiltonians of the systems and obtain their eigenenergies and eigenvectors. In Sec. III, we briefly describe the different measures of quantum correlations and state dynamical quantities. Following that, in Sec. IV, we compute the state dynamical functions and quantum correlation measures and give analytical expressions for all quantities for different initial states. Subsequently, we discuss the results in the later part of the section. Finally, in Section V, we summarise and conclude the results.
@@ II The Model Hamiltonians and methods
We are interested in exploring quantum correlations and entanglement signatures of MBSs. For this purpose, we consider a topological superconductor hosting MBSs and couple them to two QDs, as shown in Fig. 1. Moreover, we contrast this topological model by replacing the nonlocal fermion, arising due to both MBSs, with a regular fermion. In what follows, we present these two models, discuss their properties, and introduce the methods used in this work to quantify entanglement measures.
@@ II.1 QDs coupled through MBSs
where $\gamma_{i}$ represents a Majorana operator, $\epsilon_{M}$ is the energy of their splitting between MBSs, $d_{j}^{\dagger}(d_{j})$ creates (annihilates) a fermionic state in the QDs, and $\lambda_{j}$ characterizes the coupling between QDs and MBSs. We note that the energy splitting $\epsilon_{M}$ reflects the Majorana nonlocality and can be approximated as [30] $\epsilon_{M}\sim e^{-L/l}$, where L is the length of the topological superconductor and $l$ is the Majorana localisation length. Thus, $\epsilon_{M}=0$ is achieved in a very long superconductor, where $L\gg l$. In realistic systems, however, a finite value of $\epsilon_{M}$ arises due to the overlap of the wavefunctions of the MBSs [70]. Here, it is important to note that MBSs at zero energy represent true Majorana modes, whereas at finite energy, they can be seen to represent trivial ABSs [7].
To access quantum correlations and entanglement, it is convenient to write Eq. (1) in terms of fermionic operators and then handle it in number operator basis states. Thus, the two Majorana operators are transformed into a nonlocal fermion by using $\gamma_{1}=i(f^{\dagger}-f)$, $\gamma_{2}=f^{\dagger}+f$, where $f$ represents the nonlocal fermion made of Majorana operators. Now, the two QDs plus the nonlocal fermion represent a three fermion system, with number operators $n_{d_{i}}=d^{\dagger}_{i}d_{i}$ for the QDs while $n_{f}=f^{\dagger}f$ for the nonlocal fermion operator. This allows us to define a three-qubit system, which will be analyzed next.
Before proceeding, we make some simplifications by setting the on-site energies of the QDs to zero, $\epsilon_{1}=\epsilon_{2}=0$. This choice places the QDs in resonance with the truly localized MBSs, facilitating the derivation of analytical expressions; as we will see, this regime is also crucial for generating quantum correlations in the system. Further, we set the coupling to the left QD to unity: $\lambda_{1}=1$, $\omega=\epsilon_{M}/2$, $\lambda_{2}=\lambda$, and $\hbar=1$. Hence, tuning $\lambda$ and $\omega$ allows us to control the properties of the entire system. We now proceed to write Eq. (1) in the number representation basis defined by |$\ket{n_{d_{1}},n_{d_{2}},n_{f}}$. In this regard, we note that, since the Hamiltonian commutes with the fermion parity operator $(-1)^{N}$, with $N$ being the total number of fermions, the odd and the even sectors do not mix and can be solved separately. We focus on the even sector with basis states |$\ket{000}$, |$\ket{011}$, |$\ket{101}$, and |$\ket{110}$. The even Hamiltonian can be then written as,
@@
where |$\ket{E_{1}}=(1,[\omega-\Delta]/[\lambda-1],[\omega-\Delta]/[\lambda-1],1)^{T}$ and T is the transpose operation. The eigenvector |$\ket{E_{2}}=\ket{E_{1}(\Delta\rightarrow-\Delta)}$, and |$\ket{E_{3}}=(-1,[-\omega+\Delta_{1}]/[\lambda+1],[\omega-\Delta_{1}]/[\lambda+1],1)^{T}$, while |$\ket{E_{4}}=\ket{E_{3}(\Delta_{1}\rightarrow-\Delta_{1})}$. We note that all the eigenvalues in Eqs. (3) have contributions from both $\lambda$ and $\omega$, while the respective eigenstates in Eq. (4) exhibit contributions from all configurations the even ket states listed above Eq. (2). This distribution of the Hamiltonian parameters in the eigenvalues and eigenvectors arises precisely from the nonlocal fermionic structure of MBSs. We note that, although the focus of the main part is on the even sector, we also discuss the odd sector in Appendix B. However, the main findings of this work are independent of the sector.
@@ II.2 QDs coupled through a normal fermion
where $\Delta_{\pm}=\omega\pm\sqrt{1+\lambda^{2}+\omega^{2}}$. Moreover, the associated un-normalized eigenstates for increasing values of $i$ in $E^{\prime}_{i}$ are given by
@@
II.3 Methods for quantifying entanglement and state dynamics
@@
where $\rho$ is the density matrix of the composite three qubit system, and ${\rm Tr}_{d^{\prime}}$ represents the partial trace operation over a subsystem such that $\rho_{d}$ describes a two-qubit system. We remind that the three-qubit system is formed by the two QDs and either the nonlocal fermion due to MBSs or a normal fermion, see Section II for more details on the models. The density matrix in Eq. (8) is obtained by using a standard approach $\rho=\ket{\psi}\bra{\psi}$ |, where |$\ket{\psi}$ corresponds to the state of the system [71]. We anticipate that, because we are interested in the dynamical properties, we use the time-evolved states of the Majorana and normal fermion systems discussed in the previous section, taking into account the different initial states in both systems. Then, all the bipartite quantum correlations can be studied by calculating the reduced density matrix $\rho_{d}$ given by Eq. (8), thus providing a starting point for exploring various entanglement measures. In particular, we will address concurrence and discord because they provide a unifying way to quantify entanglement and quantum correlations [72, 73]. To complement these quantities, we will also address the state dynamics by employing the return probability and the entanglement dynamics. To make this work self contained, in what follows, we briefly describe the fundamental aspects of these quantities.
@@ II.3.1 Concurrence
$C=\max(\xi_{1}-\xi_{2}-\xi_{3}-\xi_{4},0)\,,$ (9)
which correspond to the maximum of the eigenvalues $\xi_{i}$, with them being the eigenvalues in decreasing order of the matrix $R=\sqrt{\sqrt{\rho_{d}}\tilde{\rho_{d}}\sqrt{\rho_{d}}}$. The matrix $\tilde{\rho_{d}}$ is defined as $\tilde{\rho_{d}}=\sigma_{y}\otimes\sigma_{y}\rho_{d}^{*}\sigma_{y}\otimes\sigma_{y}$, where $\rho_{d}^{*}$ is the complex conjugate reduced density matrix $\rho_{d}$ given by Eq. (8). In our systems, the reduced density matrices in the basis states of |$\ket{00}$, |$\ket{01}$, |$\ket{10}$, and |$\ket{11}$ acquire the following form
@@
$C=2\max(|x|-\sqrt{uv},|y|-\sqrt{w_{1}w_{2}},0)\,,$ (11)
where the maximum of the three terms is obtained by computing the eigenvalues of the matrix $R$ given just below Eq. (9), resulting in two finite entries in Eq. (11). Before going further, a few additional comments are warranted at this stage. The concurrence given by Eq. (11) remains finite if the off-diagonal correlation functions ($x,y$) dominate over their diagonal counterparts ($\sqrt{uv},\sqrt{w_{1}w_{2}}$); otherwise, $C$ goes to zero. A zero concurrence $C=0$ signals the appearance of an unentangled state of the system, but, importantly, this does not imply that the total quantum correlations in the system are completely lost or absent [76]. The approach discussed here will be used later to obtain concurrence and investigate entanglement in a bipartite system.
@@ II.3.2 Quantum discord
To quantify quantum correlations beyond entanglement, we focus on the quantum discord because it measures purely quantum correlations in a bipartite system. The quantum discord is of particular relevance when exploring the dynamics of the system, which, although acquiring a non-entangled state at certain times, may still exhibit finite quantum correlations [77, 78, 79, 80, 81, 82, 83, 84]. The quantum discord is a bipartite correlation measure, and, in this sense, it is similar to concurrence. It is defined by the difference between two classically equivalent expressions of quantum mutual information [85], which generalizes the classical mutual information to quantify the correlation between two subsystems of a bipartite system. Classically, mutual information can be defined in two equivalent ways: one using the entropies of subsystems and the other using conditional entropies [85]. In the quantum regime, the mutual information is characterized by the conditional entropy and involves measurements on one part of the system; the conditional entropy fundamentally differs from its classical counterpart because it depends on the chosen eigenbasis states. This creates a difference between the two expressions for mutual information in the quantum regime, which is quantified by quantum discord. It is thus useful to first introduce quantum mutual information, which, for subsystems A and B, is characterized by [85]
@@
where $S(\rho)$ is the von Neumann entropy associated to $\rho$ and calculated as $S(\rho)=-Tr(\rho\log_{2}\rho)$, while $C_{\theta,\phi}(\rho_{A|B})$ is the conditional entropy of A given the state of B. Moreover, $\rho_{A(B)}$ is the reduced density matrix of A(B), $\rho_{AB}$ is the composite reduced density matrix of the subsystem AB, and $\rho_{A|B}$ is the conditional density matrix of A conditioned on the measurement basis of B. Furthermore, the conditional entropy is obtained as [85]
@@
where $\theta$ and $\phi$ are angles that parametrize the measurement basis of B, $\{B_{\tilde{\kappa}}\}=\{\ket{\tilde{\kappa}}\bra{\tilde{\kappa}}\}$ | }, $\tilde{\kappa}=\{\tilde{0},\tilde{1}\}$ is a complete set of projection operators corresponding to local measurements on B, and min indicates that a minimization operation with respect to $\{B_{\tilde{\kappa}}\}$ is carried out to find $C_{\theta,\phi}(\rho_{A|B})$. Note that this minimization process involves finding the values of $\theta$ and $\phi$. Also, $p_{\tilde{\kappa}}$ on the right hand side of Eq. (13) denotes the probability of measurement outcome $\tilde{\kappa}$ and is defined as $p_{\tilde{\kappa}}=\text{Tr}[(\sigma_{0}\otimes B_{\tilde{\kappa}})\rho_{AB}(\sigma_{0}\otimes B_{\tilde{\kappa}})]$, with $\sigma_{0}$ representing the identity matrix for A. Moreover, $\rho_{\tilde{\kappa}}=(\sigma_{0}\otimes B_{\tilde{\kappa}})\rho_{AB}(\sigma_{0}\otimes B_{\tilde{\kappa}})/p_{\tilde{\kappa}}$, which represents that it is the density matrix $\rho_{AB}$ conditioned to the measurement outcomes $\tilde{\kappa}$ and weighted over different outcomes of marginal conditional entropies $S(\rho_{A|B_{\tilde{\kappa}}})$. The term $\rho_{A|B_{\tilde{\kappa}}}$ represents the conditional density matrices in the $|\tilde{\kappa}\rangle$ basis. In terms of all the above considerations, the expression for quantum discord is [85]
$D_{AB}={\rm min_{\{B_{\tilde{\kappa}}\}}}[I(\rho_{AB})-J(\rho_{AB})]\,,$ (14)
where $I(\rho_{AB})$ and $J(\rho_{AB})$ are given by Eq. (12). The minimization of the difference of two expressions over $\theta$ and $\phi$ in above equation defines quantum discord. Alternatively, it implies the computation of minimized conditional entropy, as defined in Eq. (13), which makes the quantum discord a quantifier of truly quantum correlation. In Appendix A, we provide step-by-step calculations of discord in different subsections for both MS and NFS. Here, with above considerations, we can write the expression for quantum discord
as a function of $\theta$ and $\phi$. From the above equation, it is evident that minimized conditional entropy contributes positively to quantifying discord. Therefore, a higher value of the minimized conditional entropy indicates stronger quantum discord between subsystems A and B. Additionally, the positive contribution from $S(\rho_{B})$ demonstrates that stronger entanglement between the two subsystems implies greater quantum discord. However, the composite entropy $S(\rho_{AB})$ consistently contributes negatively to the discord.